OpenStudy (anonymous):

A gardener is installing a circular garden. The circumference of the circle 132 feet. What radius should he use to layout the circle if he uses =pi 22/7 The radius is___feet.

OpenStudy (mayankdevnani):

circumference=2*pi*r 132=2*22/7*r 66=22/7*r r=21cm

OpenStudy (mayankdevnani):

@Claire4christ

OpenStudy (anonymous):

how did you get 66

OpenStudy (mayankdevnani):

by taking 2 into l.h.s and then divide 132/2=66

OpenStudy (mayankdevnani):

ok

OpenStudy (anonymous):

im not getting thesame answers explain

OpenStudy (mayankdevnani):